OpenStudy (anonymous):

find the solution of the system of equations 2x-4y=32 2x-8y=48

jimthompson5910 (jim_thompson5910):

subtract equation 2 from equation 1 to get 2x-4y=32 2x-8y=48 ---------- 0x + 4y = -16 so 4y = -16 y = ???

jimthompson5910 (jim_thompson5910):

once you determine the value of y, you can use it to find the value of x

OpenStudy (anonymous):

come on melissa....you can do this....jim pretty much gave you most of the problem

OpenStudy (anonymous):

you just have to divide everything by 4 and you get your answer

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

yes(:

jimthompson5910 (jim_thompson5910):

now that you know that y = -4, you can use this to find x so pick on any equation, plug it in and solve for x I'll use the first equation 2x-4y=32 2x-4(-4)=32 2x + 16 = 32 ... ... ... x = ???

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

8,-4

jimthompson5910 (jim_thompson5910):

(8,-4), yep

OpenStudy (anonymous):

way to go melissa...good explaining jim

jimthompson5910 (jim_thompson5910):

thanks kelliegirl33

OpenStudy (anonymous):

thanks for the help

jimthompson5910 (jim_thompson5910):

you're welcome